OpenStudy (anonymous):

The perimeter of a rectangle is 96 ft. The ratio of its length to its width is 7 : 5. What are the dimensions of the rectangle? A. 30 ft by 18 ft B. 28 ft by 20 ft C. 24 ft by 24 ft D. 22 ft by 26 ft

OpenStudy (cwrw238):

convert ratios to fractions length = 7/12 of half perimeter width = 5/12 of half parameter

OpenStudy (cwrw238):

half perimeter is of course 0.5 * 96 ft